# A unified framework for geometry-independent operator learning in cardiac electrophysiology simulations

### v2

@@ Abstract
Learning biophysically accurate solution operators for cardiac electrophysiology is fundamentally challenged by geometric variability across patient-specific heart anatomies. Most existing neural operator approaches are limited to structured or weakly deformed domains, restricting their applicability to realistic atrial and ventricular geometries. Here, we introduce a unified operator-learning framework that projects inputs and outputs onto a standardised anatomical coordinate system, decoupling electrophysiological dynamics from mesh topology. This formulation enables geometry-independent learning while preserving physiologically meaningful spatial organisation, and allows predictions to be interpolated back onto patient-specific geometries for anatomical interpretation.
To support large-scale training within the framework, we develop a GPU-accelerated electrophysiology solver and generate over 300,000 high-fidelity simulations across diverse patient-specific left atrial geometries with varied pacing and conduction properties. Within this anatomical coordinate domain, we design a neural operator to predict full-field local activation time maps, achieving a mean absolute error of 5.1 ms and an inference time of 0.12 ms per sample, outperforming existing operator learning and convolutional baselines. We further validate the framework on ventricular geometries, demonstrating robust generalisation beyond the atrial setting. Together, this framework establishes a scalable foundation for fast, geometry-invariant cardiac electrophysiology modelling, with potential relevance for real-time and population-scale clinical workflows.
@@ 1 Introduction
Computational models of cardiac electrophysiology provide a principled framework for studying electrical wave propagation, investigating arrhythmia mechanisms, and informing patient-specific therapeutic interventions [niederer2019computational, trayanova2024computational, prakosa2018personalized]. A central application of these models is the analysis of local activation time (LAT) maps, which characterise the spatiotemporal evolution of electrical activation across the myocardium [williams2018local, jaffery2024review, karoui2021cardiac]. Generating and analysing such LAT maps typically requires large ensembles of high-fidelity finite element method (FEM) simulations on anatomically detailed, patient-specific geometries, involving the numerical solution of coupled partial differential equations governing cardiac electrophysiology [niederer2011verification, opencarp].
Atrial fibrillation (AF) is a primary clinical setting in which such patient-specific LAT analyses are routinely required [sahli2018interpreting]. As the most common sustained cardiac arrhythmia, AF affects approximately 46.3 million individuals worldwide [kornej2020epidemiology] and is a leading cause of stroke [cheng2024global], heart failure [li2025global], and sudden cardiac death [lai2025causes]. Computational modelling has therefore become an important tool in AF research and clinical studies, supporting mechanistic investigation of arrhythmia drivers, evaluation of ablation strategies, and development of personalised treatment planning [heijman2021computational, boyle2019computationally, roney2021time, krogh2012effects]. However, the substantial computational cost of FEM-based workflows, often requiring hours to days per simulation, limits their applicability in settings that demand rapid or real-time decision-making [niederer2011simulating, boyle2019computationally].
Recent advances in machine learning have opened new avenues for accelerating cardiac electrophysiology simulations. Neural operators, in particular, aim to learn mappings between infinite-dimensional function spaces [kovachki2023neural], enabling solution fields to be predicted for unseen inputs, discretisations, or geometries without retraining. Architectures such as DeepONet [don] and the Fourier Neural Operator (FNO) [fno] have demonstrated strong performance on parametric partial differential equation problems. Despite this promise, most existing neural operator frameworks are developed for structured or weakly deformed computational domains, which limits their applicability to the highly irregular, patient-specific geometries encountered in cardiac modelling [geofno, gino]. In addition, training robust and generalisable operators requires large, high-quality simulation datasets, which remain expensive to generate using conventional electrophysiology solvers [ahmad2024diffeomorphic, trayanova2021machine].
Several strategies have been proposed to address anatomical variability in operator learning. Diffeomorphic mapping approaches [zhao2025diffeomorphism], including Diffeomorphic Mapping Operator Learning (DIMON) [yin2024scalable], map individual geometries to a shared reference domain, enabling learning in a common representation while encoding geometry as auxiliary information. While effective for relatively simple geometries, such as the semi-spherical left ventricle, these approaches become challenging for anatomies with complex topology, such as the multi-holed left atrium, where constructing globally consistent mappings is non-trivial. Related methods, such as Geo-FNO [geofno], deform irregular geometries to structured domains to enable spectral learning, but the accuracy and stability of such mappings may deteriorate for highly intricate anatomical structures.
@@
Here, we propose a framework that reformulates neural operator learning in a coordinate domain intrinsic to cardiac anatomy rather than in the space of mesh discretisations. By expressing model inputs and outputs in a shared anatomical parameterisation, electrophysiological patterns can be represented consistently across patients despite substantial geometric and topological variability. Predictions obtained in this coordinate space can subsequently be mapped back onto the original patient-specific geometry, enabling anatomically faithful reconstruction and direct clinical interpretation. In this work, we adopt the Universal Atrial Coordinates (UAC) [uac] and Universal Ventricular Coordinates (UVC) [uvc] systems to define such intrinsic representations for atrial and ventricular geometries, respectively. These coordinate systems preserve physiologically meaningful spatial organisation while providing a unified domain across diverse anatomies, enabling neural operators to learn spatially coherent solution mappings without sensitivity to irregular mesh connectivity or topological inconsistencies.
To support operator training within this framework, we further develop a GPU-accelerated electrophysiology simulation pipeline that enables efficient generation of high-fidelity LAT maps at scale across diverse patient-specific geometries. By integrating high-performance simulation with neural operator learning in anatomical coordinate space, our approach addresses both the computational cost of cardiac electrophysiology modelling and the geometric variability inherent in patient-specific anatomies, providing a scalable pathway toward rapid, physiology-informed inference in atrial fibrillation.
@@ 2 Results
2.1 Large-scale EP simulations on GPUs
We generated a comprehensive set of biophysically detailed EP simulations using a custom GPU-accelerated FEM solver applied to two independent cohorts of AF patients (Cohort A: $n=100$; Cohort B: $n=47$). This process constructed two large-scale datasets, Dataset A and Dataset B, respectively (see Methods). To ensure robustness across diverse anatomical and physiological conditions, for each of the $147$ patient-specific LA, we simulated LAT maps from seven distinct pacing sites across the LA. At each site, we sampled $300$ combinations of the key conduction parameters, the longitudinal conductivity ($\sigma_{l}$) and the anisotropy ratio ($\sigma_{l}/\sigma_{t}$), using Latin hypercube sampling [mckay2000comparison] to cover a wide range of clinically relevant tissue properties. This yields a total of 308,700 simulated LAT maps, establishing one of the largest computational datasets of its kind for training neural operators.
Figure 1: Construction and characterisation of large-scale electrophysiology datasets generated from biophysically detailed simulations. a, Example local activation time (LAT) map from a 3D left atrium (LA) electrophysiology simulation with coronary sinus pacing, together with its projection onto the standardised Universal Atrial Coordinate (UAC) domain. b, UMAP embedding of all UAC-projected LAT maps from Datasets A and B, showing seven pacing-site clusters per cohort and systematic displacement between corresponding clusters, indicating a pronounced domain shift between the two datasets. c, Distribution of maximum LAT across pacing locations in Dataset A, demonstrating linear dependence on LA surface area and sensitivity to conductivity. d, Corresponding analysis for Dataset B, showing the same physiological trends despite cohort-level differences.
@@ 2.1 Large-scale EP simulations on GPUs
To assess potential cohort-level differences, we performed dimensionality reduction on the full set of UAC-projected LAT maps. High-dimensional LAT fields were first denoised using principal component analysis (PCA), and the resulting principal components were subsequently embedded using Uniform Manifold Approximation and Projection (UMAP) [healy2024uniform]. The 2D embedding (Fig. 1b) reveals a clear separation between the two datasets. Within each cohort, seven well-defined clusters emerge, each corresponding to one of the pacing sites. Clusters associated with the same pacing location in Datasets A and B do not overlap but instead form displaced pairs in the latent space. This consistent displacement provides visual evidence of a domain shift between the cohorts, arising not from physiological differences, but from systematic variations in anatomical distributions, acquisition protocols, or preprocessing pipelines. This domain shift has a measurable impact on cross-domain generalisation performance. The methodology used to assess this effect is detailed in Methods.
@@
2.2 Input feature sensitivity analysis
@@
We quantified the contribution of individual anatomical input modalities by systematically removing spatial coordinates, fibre orientation, and surface area features, retraining the model under each configuration. Model performance was evaluated using mean absolute error (MAE), which captures numerical accuracy in milliseconds, and the structural similarity index measure (SSIM) [wang2004image], which assesses spatial fidelity of the predicted activation maps. All results are reported using five-fold cross-validation. Unless otherwise stated, experiments were conducted on Dataset A, with cross-domain evaluations involving both Datasets A and B.
@@
In contrast, removing fibre orientation features unexpectedly improves predictive performance across both metrics. This observation indicates that, in the present formulation, fibre orientation does not provide an additional informative signal beyond that already captured by geometric coordinates and global anatomical features. Based on this result, fibre orientation inputs were excluded from all subsequent experiments. Explanations for this effect are discussed further below.
2.3 Evaluating the effect of regularisation
@@
2.4 Benchmarking model performance
@@
Beyond accuracy, we evaluated computational efficiency to assess suitability for real-time applications. On a modern NVIDIA GPU, the model achieved an average inference time of approximately $0.12$ ms per sample. This represents orders of magnitude acceleration relative to conventional finite-element electrophysiology solvers, enabling near real-time LAT prediction across diverse anatomies and pacing configurations.
2.5 Model generalisation across pacing sites
@@
2.6 The challenge of latent data heterogeneity
@@
Integrating heterogeneous datasets during training mitigates these domain-specific biases. Training on the combined datasets $(A+B)$ encourages the model to focus on consistent anatomical features, ignoring spurious correlations tied to individual acquisition or annotation pipelines. Across both datasets, MAE decreased, and SSIM increased in comparison with the internal performance, demonstrating that exposure to independent cohorts acts as a form of regularisation. For completeness, we also evaluated the model trained and tested on the combined datasets $(A+B)\rightarrow(A+B)$, obtaining $(5.77\pm 0.16)$ for MAE and $(0.9734\pm 0.003)$ for SSIM.
2.7 Interpolation to patient-specific geometry
@@
Model performance was evaluated using a previously proposed benchmark task in which LAT is predicted from one of seven pacing locations alone [yin2024scalable]. Our experiments were conducted on Cohort A using 100 left atrial geometries, of which 20 were held out for testing. In addition, 1,006 left ventricular (LV) meshes were obtained from a publicly available dataset [lvdata], with 106 meshes reserved for evaluation. Model inputs and outputs were represented on a $100\times 100$ grid for LAs and a $50\times 50\times 50$ volumetric grid for LVs.
On test data, the model achieves mean absolute errors of 5–6 ms for LA geometries, relative to maximum activation times of up to 285 ms. For LV geometries, errors are consistently lower, typically ranging between 2 and 4 ms for maximum activation times of up to 350 ms. These ventricular error levels improve upon those reported in prior geometry-independent operator learning approaches, where mean absolute errors of approximately 5–6 ms were observed for LV meshes [yin2024scalable]. In both atrial and ventricular settings, errors remain consistent across pacing locations, indicating stable performance independent of stimulation site. Spatially, elevated errors in the LA are primarily localised to the atrial appendage and pulmonary vein openings, regions characterised by complex geometry and boundary effects. For the LV, larger errors are observed near the basal openings, corresponding to truncation at the valve plane, and along the interface with the right ventricle, reflecting geometric discontinuities introduced during extraction from bi-ventricular meshes.
@@ 3.1 Geometry-independent neural operator learning
Our approach demonstrates that a well-designed coordinate system is an efficient tool for bridging the gap between irregular, real-world data and the structured-domain requirements of neural networks, allowing a single neural operator to learn a geometry-independent mapping. By leveraging the universal coordinate system, we established a canonical representation that enabled neural operators to robustly capture the complex, nonlinear physics of cardiac conduction [hao2023gnot, li2023transformer, shih2025transformers, ovadia2024vito]. The resulting prediction accuracy, with mean absolute errors on the order of 5 ms, lies within the typical measurement uncertainty and physiological variability reported for clinical electroanatomic mapping systems [gaeta2021high] (typically on the order of 5–10 ms [narayan2024advanced]). This demonstrates that our model’s predictions are not only numerically accurate but also fall within a clinically acceptable margin.
@@
A central premise of this work is that effective neural operator learning in cardiac electrophysiology depends on access to large, diverse, and high-quality training data. To this end, we developed a GPU-accelerated finite element electrophysiology solver that enables efficient generation of high-fidelity simulations across a wide range of patient-specific left atrial anatomies and electrophysiological conditions. This data generation pipeline addresses a longstanding computational bottleneck in cardiac modelling and provides a scalable foundation for learning generalisable electrophysiological structure.
@@
Beyond supporting our framework, this large, curated dataset provides a foundation for future data-driven studies in cardiac electrophysiology. Its scale, diversity, and consistency enable predictive models to generalise across anatomies and protocols, a capability previously limited by data scarcity. For context, G-FuNK [loeffler2024graph] was trained on 25 left atrial geometries, while DIMON [yin2024scalable] leveraged 1,006 ventricular geometries simulated at multiple pacing sites. In contrast, our simulations preserve these locations while also varying tissue conductivity, allowing neural operators to capture electrophysiological responses across diverse stimulation patterns and heterogeneous, patient-specific left atrial anatomies. Overall, our dataset spans 308,700 simulations, offering a far richer foundation for training and benchmarking state-of-the-art neural operators.
The observation that removing fibre orientation features improved model performance was initially counterintuitive. Our analysis suggests two primary reasons for this finding. First, the current representation of fibre orientation may contain inherent noise or inaccuracies that interfere with the model’s ability to learn the true underlying electrophysiological dynamics. Second, the model appears to be leveraging the spatial and geometric priors provided by the coordinates and surface area features more effectively. These geometric inputs may implicitly contain sufficient information about the overall anatomical structure, making the explicit fibre orientation data redundant or even counterproductive in its current form.
3.2 Comparative model analysis
Our model achieves the best performance, which can be attributed to two key factors. First, the self-attention mechanism enables global context modelling [vaswani2017attention], allowing the network to capture long-range dependencies such as conduction wavefront spread from distant pacing sites. Second, by using convolutional patch embeddings and transposed convolution decoding, ViT maintains strong spatial localisation [wang2022convolutional], which is critical for resolving the high-frequency components of activation fronts. This balance between global awareness and local detail explains ViT’s superior pointwise accuracy and its strong preservation of spatial gradients.
We benchmarked our approach against several state-of-the-art neural operators and two different CNN architectures. Our model’s performance surpassed these baselines for reasons rooted in their fundamental design. For instance, DeepONet, which in this work employs our model for its branch or trunk networks, leverages the same global context modelling advantages as our model. This enables DeepONet to perform reasonably well, particularly in SSIM and $H^{1}$ loss, by effectively capturing smooth functional mappings and spatial dependencies. However, despite incorporating ViT components, DeepONet’s architecture focuses on decomposing input-output relationships into branch and trunk networks, which may limit its end-to-end spatial locality learning compared to the full ViT model. Consequently, its MAE performance lags behind ViT, reflecting some difficulty in the precise localisation of rapid feature changes critical for accurate cardiac wavefront propagation. In contrast, FNOs are designed to learn global mappings in function space using spectral convolutions. While FNOs can capture global trends efficiently, their reliance on low-frequency Fourier modes makes them ill-suited for capturing sharp transitions and local gradient variations common in LAT maps, particularly near pacing initiation sites. This is reflected in their relatively poor $H^{1}$ loss and MAE, despite modest SSIM performance. WNOs attempt to address this by incorporating multi-resolution representations, theoretically improving sensitivity to localised patterns. However, WNO’s performance degrades substantially in this setting, likely due to instability in wavelet decomposition on structured but irregular biomedical domains and sensitivity to small-scale alignment errors, leading to high variance in $H^{1}$ loss.
Among the convolutional baselines, U-Net delivers strong SSIM, indicating that it generates visually plausible LAT maps. Its hierarchical multi-scale architecture with skip connections allows for reconstruction of structural details. However, convolutional models like U-Net and ResNet-style CNN architectures are inherently local in receptive field, making it difficult to capture global pacing effects or subtle long-range interactions without very deep networks or large kernels [luo2016understanding].

### v3

@@ Abstract
Learning neural operators on heterogeneous and irregular geometries remains a fundamental challenge, as existing approaches typically rely on structured discretisations or explicit mappings to a shared reference domain. We propose a unified framework for geometry-independent operator learning that reformulates the learning problem in an intrinsic coordinate space defined on the underlying manifold. By expressing both inputs and outputs in this shared coordinate domain, the framework decouples operator learning from mesh discretisation and geometric variability, while preserving meaningful spatial organisation and enabling faithful reconstruction on the original geometry.
We demonstrate the framework on cardiac electrophysiology, a particularly challenging setting due to extreme anatomical variability across heart geometries. Leveraging a GPU-accelerated simulation pipeline, we generate large-scale datasets of high-fidelity electrophysiology simulations across diverse patient-specific anatomies and train customised neural operators to predict full-field local activation time maps. The proposed approach outperforms established neural operators on both atrial and ventricular geometries. Beyond cardiac electrophysiology, we further show that the same representation enables operator learning in cardiac biomechanics, a distinct problem involving volumetric deformation, highlighting the generality of the proposed framework. Together, these results establish intrinsic coordinate representations as a principled and extensible pathway for neural operator learning on complex physical systems characterised by heterogeneous geometry.
@@ 1 Introduction
Learning solution operators for partial differential equations on complex geometries remains a fundamental challenge across many areas of computational science. Neural operators offer a promising paradigm by learning mappings between functional spaces, enabling solution fields to be predicted for unseen parameters, discretisations, or geometries without retraining [kovachki2023neural]. Architectures such as DeepONet [don] and the Fourier Neural Operator (FNO) [fno] have shown strong performance on parametric PDE problems, particularly on structured or weakly deformed domains. However, extending these approaches to heterogeneous, irregular geometries remains difficult, as most neural operator formulations implicitly rely on consistent spatial discretisations or globally aligned coordinate systems [geofno, gino].
A central obstacle is that real-world physical systems are often defined on manifolds with substantial variability in shape, topology, and mesh representation. When learning is performed directly in the space of mesh discretisations, geometric differences become entangled with the underlying physical solution, hindering generalisation across geometries and limiting the applicability of neural operators to patient-specific or instance-specific domains. Existing approaches address this challenge by deforming irregular geometries to reference domains or by encoding geometry as auxiliary input [yin2024scalable, zhao2025diffeomorphism]. While effective for relatively simple shapes, such strategies can become unstable or inaccurate for domains with complex topology, multiple boundaries, or substantial variability across instances.
Here, we propose a general framework for geometry-independent neural operator learning that reformulates the learning problem in a coordinate domain intrinsic to the underlying manifold, rather than in the space of mesh discretisations. By expressing model inputs and outputs in a shared intrinsic parameterisation, spatial solution fields can be represented consistently across heterogeneous geometries despite substantial variability in shape, topology, and discretisation. Learning is performed entirely in this intrinsic coordinate domain, while predictions are subsequently mapped back to the original physical geometry for faithful reconstruction and interpretation. Importantly, the framework itself only assumes the existence of a smooth intrinsic coordinate chart on the manifold, making it broadly applicable beyond any specific physical system.
@@
Cardiac electrophysiology serves as a challenging and clinically relevant testbed for this framework, due to extreme anatomical variability, complex topology, and the need for high spatial fidelity. Computational models of cardiac electrophysiology are widely used to study electrical wave propagation, investigate arrhythmia mechanisms, and inform patient-specific therapeutic interventions [niederer2019computational, trayanova2024computational, prakosa2018personalized]. A key output of such models is the local activation time (LAT) map, which characterises the spatiotemporal evolution of electrical activation across the myocardium [williams2018local, jaffery2024review, karoui2021cardiac]. Generating LAT maps typically requires large ensembles of high-fidelity finite element simulations on anatomically detailed, patient-specific meshes, making conventional workflows computationally expensive and difficult to scale [niederer2011verification, opencarp].
@@
In this work, we adopt the Universal Atrial Coordinates (UAC) [uac] and Universal Ventricular Coordinates (UVC) [uvc] as concrete instantiations of intrinsic coordinate representations for atrial and ventricular geometries, respectively. These coordinate systems preserve physiologically meaningful spatial organisation while providing a unified domain across diverse anatomies, enabling neural operators to learn spatially coherent solution mappings without sensitivity to irregular mesh connectivity or topological inconsistencies. To support operator learning at scale, we further introduce an efficient GPU-accelerated electrophysiology simulation pipeline that enables the generation of large ensembles of biophysically detailed LAT maps across diverse patient-specific geometries.
By integrating intrinsic coordinate representations with high-performance simulation and neural operator learning, the proposed framework addresses both the geometric heterogeneity and the computational cost that have limited prior approaches. While cardiac electrophysiology serves as a challenging and clinically relevant demonstration, the framework is broadly applicable to other physical systems in which intrinsic coordinates are available, illustrating a general pathway toward geometry-independent operator learning on complex manifolds.
@@ 2 Results
2.1 Large-scale electrophysiology datasets in intrinsic coordinate space
We first assess whether an intrinsic coordinate representation yields a consistent learning domain across heterogeneous geometries, using LAT map prediction in cardiac electrophysiology as a representative operator learning task.
We generated a comprehensive set of biophysically detailed EP simulations using a custom GPU-accelerated FEM solver applied to two independent cohorts of AF patients (Cohort A: $n=100$; Cohort B: $n=47$). This process constructed two large-scale datasets, Dataset A and Dataset B, respectively (see Methods). To ensure robustness across diverse anatomical and physiological conditions, for each of the $147$ patient-specific left atriums (LAs), we simulated LAT maps from seven distinct pacing sites across the LA. At each site, we sampled $300$ combinations of the key conduction parameters, the longitudinal conductivity ($\sigma_{l}$) and the anisotropy ratio ($\sigma_{l}/\sigma_{t}$), using Latin Hypercube Sampling (LHS) [mckay2000comparison] to cover a wide range of clinically relevant tissue properties. This yields a total of 308,700 simulated LAT maps, establishing one of the largest computational datasets of its kind for training neural operators.
Figure 1: Construction and characterisation of large-scale electrophysiology datasets generated from biophysically detailed simulations. a, Example local activation time (LAT) map from a 3D left atrium (LA) electrophysiology simulation with coronary sinus pacing, together with its projection onto the standardised Universal Atrial Coordinate (UAC) domain. b, UMAP embedding of all UAC-projected LAT maps from Datasets A and B, showing seven clusters per cohort corresponding to pacing sites and systematic displacement between corresponding clusters, indicating a pronounced domain shift between the two datasets. c, Distribution of maximum LAT across pacing locations in Dataset A, demonstrating linear dependence on LA surface area and sensitivity to conductivity. d, Corresponding analysis for Dataset B, showing the same physiological trends despite cohort-level differences.
@@ 2.1 Large-scale electrophysiology datasets in intrinsic coordinate space
To characterise variability within this intrinsic representation, we performed dimensionality reduction on the full set of UAC-projected LAT maps. High-dimensional LAT fields were first denoised using principal component analysis (PCA), and the resulting principal components were subsequently embedded using Uniform Manifold Approximation and Projection (UMAP) [healy2024uniform]. The 2D embedding (Fig. 1b) reveals a clear separation between the two datasets. Within each cohort, seven well-defined clusters emerge, each corresponding to one of the pacing sites. Clusters associated with the same pacing location in Datasets A and B do not overlap but instead form displaced pairs in the latent space. This consistent displacement provides visual evidence of a domain shift between the cohorts, arising not from physiological differences, but from systematic variations in anatomical distributions, acquisition protocols, or preprocessing pipelines. This domain shift has a measurable impact on cross-domain generalisation performance. The methodology used to assess this effect is detailed in Methods.
@@
2.2 Sensitivity to intrinsic representation and input features
@@
We quantified the contribution of individual anatomical input modalities by removing spatial coordinates, fibre orientation, and surface area features, retraining the model under each configuration. Model performance was evaluated using mean absolute error (MAE), which captures numerical accuracy in milliseconds, and the structural similarity index measure (SSIM) [wang2004image], which assesses spatial fidelity of the predicted activation maps. All results are reported using five-fold cross-validation. Unless otherwise stated, experiments were conducted on Dataset A, with cross-domain evaluations involving both Datasets A and B.
@@
In contrast, removing fibre orientation features unexpectedly improves predictive performance across both metrics. This observation indicates that, in the present formulation, fibre orientation does not provide an informative signal beyond that already captured by geometric coordinates and global anatomical features. Consequently, fibre orientation inputs were excluded from all subsequent experiments. This finding does not imply that fibre orientation is unimportant for cardiac electrophysiology, but rather that its current atlas-based representation, when projected into the intrinsic coordinate domain, does not contribute additional predictive value for this task.
2.3 Effect of spatial regularisation on operator learning
@@
2.4 Comparison with existing neural operator architectures
@@
Beyond accuracy, we evaluated computational efficiency to assess suitability for real-time applications. On a modern NVIDIA GPU, the model achieved an average inference time of approximately $0.12$ ms per sample. This represents orders of magnitude acceleration relative to conventional FEM electrophysiology solvers, enabling near real-time LAT prediction across diverse anatomies and pacing configurations.
2.5 Generalisation across stimulation conditions
@@
2.6 Generalisation under latent domain shifts
@@
Integrating heterogeneous datasets during training mitigates these domain-specific biases. Training on the combined datasets $(A+B)$ encourages the model to focus on consistent anatomical features, ignoring spurious correlations tied to individual acquisition or annotation pipelines. Across both datasets, MAE decreased, and SSIM increased in comparison with the internal performance, demonstrating that exposure to independent cohorts acts as a form of regularisation and that geometry-independent representations alone are insufficient for translational robustness unless coupled with heterogeneous, multi-domain training. For completeness, we also evaluated the model trained and tested on the combined datasets $(A+B)\rightarrow(A+B)$, obtaining $(5.77\pm 0.16)$ for MAE and $(0.9734\pm 0.003)$ for SSIM.
2.7 Reconstruction on original geometries
@@
Model performance was evaluated using a previously proposed benchmark task in which LAT is predicted from one of seven pacing locations alone [yin2024scalable]. Our experiments were conducted on Cohort A using 100 LA geometries, of which 20 were held out for testing. In addition, 1,006 LV meshes were obtained from a publicly available dataset [lvdata], with 106 meshes reserved for evaluation. Model inputs and outputs were represented on a $100\times 100$ grid for LAs and a $50\times 50\times 50$ volumetric grid for LVs.
On test data, the model achieves mean absolute errors of 5–6 ms for LA geometries, relative to maximum activation times of up to 285 ms. For LV geometries, errors are consistently lower, typically ranging between 2 and 4 ms for maximum activation times of up to 350 ms. These ventricular error levels improve upon those reported in prior geometry-independent operator learning approaches, where mean absolute errors of approximately 5–6 ms were observed for LV meshes [yin2024scalable]. The comparison is based on matched dataset scale, task definition, and activation time range, rather than a direct reimplementation, which is not feasible because the prior work relies on a simplified LV representation that is not compatible with the extracted LV meshes considered here.
In both atrial and ventricular settings, errors remain consistent across pacing locations, indicating stable performance independent of stimulation site. Spatially, elevated errors in the LA are primarily localised to the atrial appendage and pulmonary vein openings, regions characterised by complex geometry and boundary effects. For the LV, larger errors are observed near the basal openings, corresponding to truncation at the valve plane, and along the interface with the right ventricle, reflecting geometric discontinuities introduced during extraction from bi-ventricular meshes.
2.8 Application to kinematic deformation modelling
To showcase that the proposed approach is not specific to cardiac electrophysiology, we apply it to a numerically distinct problem: kinematic deformation of non-linear elastic solids under load. This task involves volumetric deformation governed by material stiffness and applied pressure, and therefore represents a fundamentally different class of physical system and boundary conditions.
We generate a dataset of three-dimensional geometries by varying two independent geometric parameters controlling global curvature and wall thickness. For each geometry, kinematic deformation simulations are performed across different combinations of material stiffness and internal pressure under fixed boundary constraints. As in the cardiac electrophysiology setting, a problem-specific intrinsic coordinate system is constructed to provide a consistent parameterisation across heterogeneous meshes, enabling all displacement fields to be represented in a shared coordinate domain.
The neural operator takes four scalar inputs describing geometry and loading conditions and predicts the full-field displacement magnitude. Performance is evaluated using ten-fold cross-validation. The average displacement magnitude in the dataset is approximately $2\,\mathrm{mm}$, with a maximum of about $7\,\mathrm{mm}$, while the model achieves a mean absolute error of $0.35\pm 0.28\,\mathrm{mm}$ across test folds. Relative to the characteristic displacement magnitudes in the dataset, this error level indicates that the learned operator captures the dominant deformation behaviour across heterogeneous geometries.
@@ 3.1 Geometry-independent neural operator learning
A key insight arising from this work is that geometry independence in operator learning can be most effectively addressed at the level of representation rather than architecture. Within this framework, the role of the intrinsic coordinate system is not to simplify the underlying physics, but to provide a canonical representation in which spatially coherent learning becomes feasible. By embedding geometric structure directly into the learning representation, neural operators can exploit both local spatial relationships and global contextual information without being confounded by irregular mesh connectivity or patient-specific discretisation artefacts. Importantly, the framework itself is agnostic to the specific physical system, requiring only the availability of a smooth intrinsic coordinate chart that is consistent across instances. Potential applications extend to include fluid–structure interaction, geophysical modelling on planetary surfaces, and biological transport processes on irregular anatomical domains.
@@
Cardiac electrophysiology provides a particularly stringent and clinically relevant testbed for this methodology, due to the pronounced variability in anatomy, topology, and boundary conditions across patients. By leveraging universal anatomical coordinate systems, we instantiate the proposed framework in this setting and demonstrate that neural operators can capture the complex, nonlinear dynamics of electrical conduction across highly heterogeneous cardiac geometries. The resulting prediction accuracy, with mean absolute errors on the order of 5 ms, lies within the typical measurement uncertainty and physiological variability reported for clinical electroanatomic mapping systems [gaeta2021high] (typically 5–10 ms [narayan2024advanced]). This indicates that the learned operators are not only numerically accurate but also operate within a clinically acceptable margin.
@@
A further premise underlying geometry-independent operator learning is the need for large, diverse, and high-quality training data. To support this, we developed a pipeline featuring a GPU-accelerated finite element electrophysiology solver that enables efficient generation of high-fidelity simulations across a wide range of patient-specific left atrial anatomies and electrophysiological conditions. This data generation capability addresses a longstanding computational bottleneck in cardiac modelling and provides a scalable foundation for learning generalisable structure within the proposed framework.
@@
Beyond enabling the present study, the resulting dataset establishes a resource for future data-driven investigations. Its scale, diversity, and internal consistency allow predictive models to generalise across anatomies and simulation protocols, a capability that has previously been constrained by data scarcity. For comparison, G-FuNK [loeffler2024graph] was trained on 25 left atrial geometries, while DIMON [yin2024scalable] leveraged 1,006 ventricular geometries simulated at multiple pacing sites, yielding 7,042 simulation instances. In contrast, our simulations preserve multiple pacing locations while also varying tissue conductivity, allowing neural operators to capture electrophysiological responses across diverse stimulation patterns and heterogeneous, patient-specific anatomies. Overall, the dataset comprises 308,700 simulations, offering a substantially richer foundation for training and benchmarking state-of-the-art neural operators.
3.2 Inductive biases for learning on heterogeneous geometries
The comparative evaluation reveals that strong performance in geometry-independent operator learning depends less on the choice of a specific neural architecture than on the inductive biases it encodes. In particular, accurate reconstruction of spatial fields on heterogeneous domains requires a balance between global contextual reasoning and precise local spatial representation. Models that emphasise only one of these aspects consistently underperform in this setting.
Our approach benefits from this balance by combining global context modelling with explicit spatial structure, enabling the network to capture long-range dependencies while preserving fine-scale localisation. This is particularly important for activation time prediction, where global wavefront propagation and sharp local gradients coexist within the same spatial field. Architectures that capture global dependencies without maintaining sufficient spatial fidelity struggle to resolve rapid local transitions, while those that prioritise local structure alone fail to model long-range interaction driven by wavefront propagation.
These trends are consistent across the evaluated baselines. Operator learning methods designed to model smooth global mappings perform well in terms of structural similarity but exhibit reduced pointwise accuracy in regions with steep spatial gradients. Conversely, convolutional architectures with strong local inductive biases generate visually coherent predictions yet remain limited in their ability to account for global pacing effects without substantial increases in depth or receptive field. Multi-resolution approaches offer partial mitigation but remain sensitive to alignment errors and domain irregularities in this context.
Together, these results indicate that geometry-independent operator learning on irregular anatomical domains places specific demands on model design. Architectures must simultaneously integrate global information and preserve spatial locality in order to achieve both numerical accuracy and structural fidelity. This requirement is largely independent of the specific application and is expected to generalise to other problems involving partial differential equations on complex geometries.